Julie is typing her essay at a rate of 300 words per hour. If on Monday, Julie typed 1000 words, how many hours does Julie spend on Tuesday typing, if by the end of the day her essay has a total of 3600 words?
Found 2 solutions by Theo, FrankM:
Answer by Theo(13342)

rate * time = quantity.
r * t = q
r = rate
t = time
q = quantity

r = 300
q = 1000
r * t = q becomes 300 * t = 1000
it took 1000/300 = 3 and 1/3 hours to type 1000 words on monday.

by the end of the day on tuesday, she had typed 3600 words.
this included the 1000 words she typed on monday.
she must have typed 2600 words on tuesday.
r * t = q becomes 300 * t = 2600.
it took 2600 / 300 = 8 and 2/3 hours to type 2600 words on tuesday.

your solution should be 8 and 2/3 hours.

she typed a total of 3600 words.
r * t = q becomes 300 * t = 3600
t = 3600 / 300 = 12 hours total.

she typed 3 and 1/3 hours on monday and she typed 8 and 2/3 hours on tuesday.
that's a total of 12 hours.
the solution that she spent 8 and 2/3 hours on tuesday is confirmed correct.





Answer by FrankM(1040)

Take it one step at a time.
She went from 1000 words to 3600 words. That's 2600 words typed on Tuesday.
2600 words * 1hr/300words = 8 2/3 hours typing.

The lesson here is how to think of the rate 300 words per hour. If I asked you how many words can she type in 4 hours, you'd say
4 hours * 300 words/hour = 1200. And if you notice, hours/hours cancel and we are left with an answer in number of words.
For our problem, we have words, and need to get to hours. We have to flip 300words/hour to 1hr/300words, which will solve the conversion properly.